Shift report 20010920 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Sep 20 2001 - 00:31:09 EDT

  • Next message: Apache: "Shift report 20010920 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Sep 20 00:31:09 2001, Shift: 16:00-24:00
    Supervisor: yklee, with: 
    Log book#9 pages 11 - 12
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine had RF-high level problem and ramping problem
      and the first beam for the shift started around 23:00.
    2. Run 4817 has 150 K trigger 1, and is still running.
       We are still working with
    	FS 8 degrees 1/5 A polarity
    	MRS 90 degrees 4B A polarity
    3. ZDC Low gain ADC Blue Right module 1 is missing.
       T3 and T5 tripped once each.
       Magnet technician still could not fix D5 remote control
       problem. He asked us to call MCR and get some one to
       replace the module powering A to D converter.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Sep 20 2001 - 00:31:30 EDT